OpenStudy (anonymous):

functions f an g defined as f(x)=2x-5 an g(x)=x[2]+3 what is f(4) gf(4) an the inverse of f (x)

OpenStudy (anonymous):

f(4)=2(4)-5 =3 gf(4)=g(3) =3(2)+3 =9 f(x)=2x-5 which is the same as y=2x-5 first step is to interchange the values x and y x=2y-5 then we make y the subject of the formula x+5=2y (x+5)/2=y

OpenStudy (anonymous):

thank you